OpenStudy (anonymous):

I am doing graphs of functions and relations; need to find the domain and range for x=(y+2)^2

OpenStudy (ericdl):

domain: [0,∞) range: (-∞,∞)

OpenStudy (anonymous):

to find the domain we make the equation y=? so we squareroot the equation to get y+2=sqrt(x) then subtract 2 y=sqrt(x)-2 do you know what a squareroot function looks like?

OpenStudy (anonymous):

this one is only moved down 2 units